OpenStudy (anonymous):

if 1.4 pounds of beef costs 3.64 dollars how much does a 1 pound costs?

OpenStudy (anonymous):

1.4-3.64 1-x x=3.64/1.4=2.6

OpenStudy (ajprincess):

Let us say it cost x dollars. 1.4->3.64 1->x 1.4/1=3.64/x x=3.64/1.4

hero (hero):

\[\frac{1.4 \text{beef}}{1 \text{beef}} = \frac{3.64 \text{dollars}}{x \space\text{dollars}}\] Cancel beef and dollars to get: \[\frac{1.4}{1} = \frac{3.64}{x}\] Cross multiply to get 1.4x = 3.64 Divide both sides by 1.4 to get x = 3.64/1.4 x = 2.6

OpenStudy (anonymous):

ty all

OpenStudy (ajprincess):

yw:)

OpenStudy (anonymous):

:-)